OpenStudy (anonymous):

The sum of three numbers is 60. The second number is two more than the first. The third number is two more than the second. Find the three numbers

OpenStudy (jim766):

if the first number is x then the 2nd number is x + 2 two more than 1st then the 3rd number is x + 2 + 2 two more than the 2nd now just add them up = 60

OpenStudy (jim766):

x + x + 2 + x + 2 + 2 = 60 solve for x

OpenStudy (uri):

(x) + (x+2) + (x+4) = 60 3x + 6 = 60 3x=60-6 3x = 54 x=54/3 x=18 therefore the second number is 20 because 18+2=20 and the third is 22 because 20+2=22 :)

OpenStudy (anonymous):

Thanks so much.

OpenStudy (anonymous):

let first number is x then second number is x+2 so the third number is x+2+2 as per question (x)+(x+2)+(x+2+2) = 60 3x+6 = 60 3x =60-6 x= 54/3 =18 so 1st number is 18 second number is 20 and third number is 22 You can practice different number properties : http://www.mathcaptain.com/number-sense/properties-of-numbers.html problems for your improvement.